Aggregator

New RTP HV Drivers - Hall A Gateway Channel Access

6 days 3 hours ago
Log Number: 4467978Logbook(s): Entry Makers: suleimanBody: 

For the old driver, we have - see linked entry:
"3. DACs of new RTP cell:
IGL0I00C1068_DAC05, IGL0I00C1068_DAC06, IGL0I00C1068_DAC07, IGL0I00C1068_DAC08, IGL0I00C1068_DAC09, IGL0I00C1068_DAC10, IGL0I00C1068_DAC11, IGL0I00C1068_DAC12."

We should keep these for the time being in case we need to go back to the old driver.

For the new drivers, we should add these PVs to the Hall A Gateway Channel Access:

RTPXPORT1Q1calc.A
RTPXPORT1Q4calc.A
RTPXPORT2Q1calc.A
RTPXPORT2Q4calc.A

RTPXPORT1Q2calc.A
RTPXPORT1Q3calc.A
RTPXPORT2Q2calc.A
RTPXPORT2Q3calc.A

suleiman

New RTP HV Driver - Alarm Handler

6 days 3 hours ago
Log Number: 4467976Logbook(s): Entry Makers: suleimanBody: 

For the new RTP HV Drivers, we want to add these signals to the alarms. We want to keep in the alarms the old RTP Driver but have it disabled, in case we have to go back to the old one.

Under LASER-INJECTOR-POCKELS CELL, we should add both "RTP 1 HV DRIVER" and "RTP 2 HV DRIVER". The two signals are "RTPXPORT1DCTRL" and "RTPXPORT2DCTRL", "1" ON, "0" OFF.

We want also to add to the Alarm the eight HVs.

suleiman

Set up to FC2

6 days 22 hours ago
Log Number: 4467969Logbook(s): Entry Makers: samaribschaumlBody: 

We initially centered through the buncher, then proceeded to fine tune dechopping, while making sure that all beams time of arrival to the booster was equal.
We checked D Laser bucket and verified 750 MHz timing on IPM1I02.

Then we checked booster phasing, and found that the 7 cell was on crest while the 2 cell had been on phase then 35.5 degrees had been subtracted. We left it as we found it at the beginning of shift after scratching our heads for far too long trying to find out which cavity should be off crest.

samari

Autolog save

6 days 22 hours ago
Log Number: 4467966Logbook(s): Tags: Body: saver ran from opsl632

Saved: samari 01-31-26 13:38:36 {A:5 B:3 C:1 E:340 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in. dump(s) in. Initial set up to FC2..
samari

HarpFitter: IHA0L03

6 days 22 hours ago
Log Number: 4467965Logbook(s): Body: 

Comments:

After putting 0L quads on design

Laser Used:

A Laser

Sigmas:

Peak 1 - x Sigma: 0.2215 Peak 2 - u Sigma: 0.3802 Peak 3 - y Sigma: 0.6360
samari

HYST AREA 7-0 Cycles Magnets for eDT

6 days 22 hours ago
Log Number: 4467964Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

Magnets contained in the following file: /tmp/eDT_CMPQ_design_snapfile_0gh0jB.snap
MQJ0L01
MQJ0L02
MQJ0L02A
MDL0L02
MQJ0L03A
MQJ0L03
MQJ0L04
MQD0L06
MQB0L07
MQB0L08
MQB0L09
MQB0L10

Total Time: 1.283 mins

samari

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Blackshaw
MCC Operators Thomas
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph